OpenStudy (australopithecus):

I'm trying to calculate the volume of y = 2-x^(2) y = x^(2) Rotated around x=1 using shells but I keep getting an integral that gives me zero

OpenStudy (australopithecus):

\[\int\limits_{-1}^{1} 2pi(x)(2-x^{2} - x^{2})dx = 0\]

OpenStudy (australopithecus):

the height being (2−x2−x2) am I incorrect that the radius is x?

OpenStudy (australopithecus):

This being the picture of the solid and shell

OpenStudy (anonymous):

yep.. your radius should be 1-x

OpenStudy (australopithecus):

why is that if I may ask?

OpenStudy (anonymous):

the distance from your rotation axis to the representative rectangle with width dx is 1- the x-coordinate. so r=1-x